# Strange metal and superconductor in the two-dimensional Yukawa-Sachdev-Ye-Kitaev model

### v3

@@ Abstract
Moreover, YSYK models also display instabilities of the strange metal to superconductivity [14, 15, 16, 17, 18], with the pairing type dependent upon the particular quantum phase transition being studied. We will examine an instability to spin-singlet pairing in a simplified model which ignores the gap variation around the Fermi surface, and will therefore only apply to the anti-nodal regions of the cuprates. We present the evolution of the electron spectral function and superfluid density for $T<T_{c}$, the superconducting critical temperature. We find that the 2d-YSYK model exhibits a number of experimentally observed trends:
@@
The 2d-YSYK Model—Several works [23, 24, 25, 26, 27, 28, 29, 30] have argued that clean quantum critical metals cannot serve as a universal model for transport in a strange metal, and that impurity-induced spatial disorder is essential. As a characteristic example, we therefore add Harris disorder to the Hertz-Millis theory of a quantum phase transition in a metal associated with an Ising-nematic order parameter $\phi$ [31]. Such disorder is provided by quenched random terms which preserve the Ising symmetry [25, 26]. Other order parameters, including those at non-zero wavevector, and Fermi-volume changing transitions without broken symmetries [32], also map to essentially the same 2d-YSYK model [8]. We define the 2d-YSYK model here as the Harris-disordered Hertz-Millis Lagrangian for $\phi$ and fermions $\psi$ with dispersion $\varepsilon({\bm{k}})$ [7, 33, 8, 28, 30, 34]:
@@
The flavor indices $i,j,\ell$ are summed over $N$ values—the equations in Fig. 1 are exact at all $T$ in the large $N$ limit when the couplings are random in flavor space. Our discussion is for the physical case $N=1$, for which Fig. 1 is applied over an intermediate $T$ regime. Imaginary time is $\tau$, we set $\hbar=1$, $s$ is the tuning parameter across the transition, $u$ is a scalar self-interaction. The operator ${\mathcal{D}}_{\bm{r}}=\partial_{x}^{2}-\partial_{y}^{2}$ is special to the Ising-nematic case, and will be set to unity for simplicity in our computations as it unimportant apart from ‘cold spots’ on the Fermi surface. In order to obtain a spin-singlet superconductor we have spinful fermions, with $\sigma$ the spin index.
$\mathcal{L}_{\rm 2d-YSYK}$ contains the two sources of disorder. One is the potential $v({\bm{r}})$ acting on the fermions. It is random in both flavor and space (in (2) and (3) we omit flavor indices for clarity, see the supplement [22] for full forms):
@@
(i) There is a significant intermediate energy regime where the bosonic and fermionic eigenmodes are spatially extended. In this regime, the physics is self-averaging, and numerical results are consistent with the large-$N$ averaged Green’s function methods which yield the exact solution of the zero-dimensional YSYK model. This is the regime we address in the present paper by standard SYK methods. The spatial disorder is needed for the singular single-particle self-energies to feed into singular transport properties [8], which does not happen for translationally invariant models [23, 24, 25, 26, 27, 28, 29, 30].
@@
Normal State Results—As we describe in the supplement [22], the self-consistent equations for the Green’s functions can be obtained by endowing the fields with flavor indices, and making the Yukawa coupling also random in flavor space. After a disorder average, and in the limit of a large number of flavors, we obtain the SYK-type equations which are indicated schematically in Fig. 1.
@@
Our numerical findings for the phase diagram are summarized in Fig. 2a, which is plotted as a function of the renormalized boson mass $M$ used to tune the system to the quantum critical point (QCP), and $T$. The phenomenology is broadly similar to that observed experimentally in a wide variety of strange metals: Above the QCP there is a quantum-critical fan, in which the resistivity has an approximately linear $T$ dependence. At low $T$ the QCP is ultimately masked by a superconducting phase, with the maximal $T_{c}$ occurring at the critical point. The d.c. transport is shown in more detail in Fig. 2b. Evidently, the logarithmic corrections to the resistivity mentioned above are relatively weak for the parameters we have considered. On the disordered side of the transition the resistivity is $T$-linear at elevated $T$ inside the quantum critical fan, before going to zero with what is well-approximated by a $T^{2}$ power-law below a certain crossover temperature. Over the entire $T$ range shown the resistivity is smaller than the quantum of resistance; the system is not a bad metal.
@@
We now consider the finite-frequency response  [^1] The related problem of optical conductivity at a clean Ising-nematic QCP has been recently analyzed in [51, 52].. The real and imaginary parts of the optical conductivity are presented in Figs. 3a and b. At a crude level, the optical conductivity can be seen to have an approximately Drude-like behavior. To better understand the more detailed structure encoded in $\sigma(\omega)$ we follow Michon et. al. [6] and parametrize $\sigma(\omega)$ via a ‘generalized’ Drude formula:
@@
Here $K$ is the optical weight and is equal to the average electronic kinetic energy (see (LABEL:Kdef) in the supplement [22]). With $K$ defined in this way, we may read off a frequency-dependent transport scattering rate $1/\tau_{\text{tr}}(\omega)$ and frequency dependent mass-enhancement parameter $m^{*}(\omega)/m$ directly from the real and imaginary parts of the optical conductivity, respectively.
@@
The optical scattering rate is shown in Fig. 3c. There is an approximately linear in frequency dependence down to $\omega\sim T$, while for frequencies $\omega\lesssim T$ the $1/\tau_{\text{tr}}$ tends to a $T$-dependent non-zero value which vanishes in the limit $T\to 0$. The larger $\omega$ values of the dimensionless ratio $1/[\omega\tau_{\rm tr}(\omega)]$ in Fig. 3c is smaller than the observed value [6], but larger values of the ratio appear at larger $g^{\prime}$, as shown in Fig. LABEL:fig:sig_g_5_insetb [22].
@@
The frequency-dependent mass enhancement parameter is shown in Fig. 3d. For the chosen parameters, the low-frequency mass enhancement does not exceed roughly 20% down to the lowest temperatures we are able to reliably access in the numerical calculations. The behavior of this “optical” mass enhancement is consistent with the mass enhancement we have inferred from the fermion self-energy [22]. At sufficiently low temperatures $m^{*}(\omega=0)/m$ is expected to diverge logarithmically with $T$ [8]. While we do not observe a pure logarithmic growth of $m^{*}/m$ in our data, the mass enhancement does continuously increase with decreasing $T$ at the QCP, suggesting a (slow) divergence as $T\to 0$.
@@
We also report the modulus and phase of the optical conductivity in Figs. 3e and f. Notably, we find the modulus of the conductivity has an apparent sub-linear power law behavior over an intermediate frequency range $|\sigma|\sim 1/\omega^{\nu^{*}}$, where $\nu^{*}\simeq 0.9$ for the chosen set of parameters. Over a similar frequency range the phase of the optical conductivity plateaus at a value $\arg\sigma\simeq\nu^{*}\pi/2$. Such behavior at intermediate frequencies has been observed in infrared conductivity measurements of cuprates [6, 5, 49]. The exponent $\nu^{*}$ is continuously tunable with parameters and, in particular, is a decreasing function of the coupling strength $g^{\prime}$ [22].
@@
Within the quantum critical fan, frequency-dependent observables are generally expected to obey $\omega/T$ scaling. In the present model, $\omega/T$ scaling for the optical conductivity is spoiled by logarithmic corrections [8, 6], the most significant such effect arising from the logarithmic divergence of the effective mass with $T$. This may be accounted for by utilizing the generalized Drude formula (4) and considering separately $\omega/T$ scaling of $1/\tau_{\text{tr}}$ and $\Delta m^{*}/m=m^{*}(\omega)/m-m^{*}(0)/m$, where the subtraction removes a contribution that is expected to violate scaling. With this approach we find reasonable scaling collapse for the optical scattering rate – see inset of Fig. 3c. For the parameters presented here, significant logarithmic corrections apparently remain for $\Delta m^{*}/m$. To the extent there is any reasonable scaling collapse, it only holds over a much narrow range of $\omega/T$ – see inset of Fig. 3d.
@@
Superconductivity Results—We now investigate the onset of superconducting pairing in our theory. The superconducting transition temperature $T_{c}$, as shown in Fig. 2a, is numerically identified by the linearized gap equation; see Appendix LABEL:sec:real_YSYK. For fixed $g^{\prime}$, as we tune away from the QCP by increasing the renormalized boson mass $M$, we find that $T_{c}$ decreases. We compare $T_{c}$ with $A$, the slope of the linear-$T$ resistivity, in the inset of Fig. 2b. The relation is monotonic, but not linear as discussed by Taillefer [19]. We note, however, that in [19] the coefficient $A$ was not extracted in the quantum-critical fan, which is how we have defined it, but rather from fitting to the low-$T$ resistivity in an “extended" critical regime – this regime has been associated with the localization of $\phi$ [9], and is not addressed by our analysis here. If we remain at criticality while changing $g^{\prime}$, then we do find a linear relationship between $A$ and $T_{c}$, as shown in Fig. LABEL:fig:Tca [22].
@@
In Fig. 4 we show the evolution of the electronic density of states at the quantum critical point as the system goes through the superconducting transition, demonstrating an apparently conventional gap opening as temperatures decreases to $T<T_{c}$ despite the fact that, in the absence of superconductivity, the fermions behave as a marginal Fermi liquid with $-\mathrm{Im}\Sigma\sim|\omega|$ at low $T$, as seen in the inset of Fig. 4.
@@
We have also computed the superfluid stiffness, which may be expressed in terms of the anomalous Matsubara Green’s function $F$ as
$\frac{\rho_{s}}{\pi e^{2}}=\frac{2}{\beta}\sum_{\omega_{n}}\int{\rm d}\epsilon~{}\rho_{\text{tr}}(\epsilon)F(\epsilon,i\omega_{n})F^{\dagger}(\epsilon,i\omega_{n}),$ (5)
where the density of states $\rho_{{\rm tr}}^{\alpha\beta}\left(\epsilon\right)\equiv\int_{\bm{k}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\alpha}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\beta}}\delta\left(\epsilon-\varepsilon_{\bm{k}}\right)$. In Fig. 5, we show the temperature dependence of $\rho_{s}$ for $g^{\prime}=5t^{3/2}$ as the system is tuned away from the QCP by varying the boson mass $M$. The stiffness reaches a finite constant value, $\rho_{s0}$, for $T\rightarrow 0$. As the system is tuned away from the QCP, $T_{c}$ decreases from its maximal value while $\rho_{s0}$ increases, as seen in Fig. 5. This result at strong coupling aligns with the superfluid density in bulk samples of the overdoped cuprate superconductors [50, 20]. When $T_{c}$ is largest at the QCP, the opening of a $\psi$ gap at $T=0$ weakens the pairing interaction mediated by $\phi$ (because the $\phi$ self energy $\Pi$ is self-consistently determined by the $\psi$ polarization, see Fig. 1), leading to a smaller superfluid stiffness at $T=0$ [17, 18]. On the other hand, in the presence of strong potential scattering $v$ this effect is weaker, and a decreasing $T_{c}$ is eventually accompanied by a decreasing $\rho_{s0}$, as shown in Fig. LABEL:fig:stiffness_s [22]. The low $T$ boson localization [9] could have a significant influence on the spatial inhomogeneity of the superfluid density, and this remains to be studied.
@@
Discussion—The spatially inhomogeneous fermion-boson coupling $g^{\prime}({\bm{r}})$ in the 2d-YSYK model in (1) is an alternative to ‘random mass’ $\delta s({\bm{r}})$ disorder in Hertz-Millis theory, in the crossover regime where the bosonic modes remain extended despite the disorder [9], as we discussed below (3). The striking similarities between the results presented here and a variety of properties measured in strange metal superconductors such as the cuprates – including static and dynamic properties across the normal and superconducting states – indicate such disorder plays a significant role in the phenomenology of these systems, making it important to determine the microscopic origin of disorder configurations that strongly affect the local position of the QCP. Finally, the crucial issue of how our results evolve as the bosonic modes localize due to disorder [9] will be addressed in future work.

### v4

@@ Abstract
Moreover, YSYK models also display instabilities of the strange metal to superconductivity [14, 15, 16, 17, 18], with the pairing type dependent upon the particular quantum phase transition being studied. We examine an instability to spin-singlet pairing in a simplified model which ignores the gap variation around the Fermi surface, and therefore only applies to the anti-nodal regions of the cuprates. We present the evolution of the electron spectral function and superfluid density for $T<T_{c}$, the superconducting critical temperature. We find that the 2d-YSYK model exhibits a number of experimentally observed trends:
@@
The 2d-YSYK Model—Several works [23, 24, 25, 26, 27, 28, 29, 30] have argued that clean quantum critical metals cannot serve as a universal model for transport in a strange metal, and that impurity-induced spatial disorder is essential. We therefore add Harris disorder to the Hertz-Millis theory of a quantum phase transition in a metal associated with an Ising-nematic order parameter $\phi$ [31]. Such disorder is provided by quenched random terms which preserve the Ising symmetry [25, 26]. Other order parameters, including those at non-zero wavevector, and Fermi-volume changing transitions without broken symmetries [32], also map to essentially the same 2d-YSYK model [8]. We define the 2d-YSYK model here as the Harris-disordered Hertz-Millis Lagrangian for $\phi$ and fermions $\psi$ with dispersion $\varepsilon({\bm{k}})$ [7, 33, 8, 28, 30, 34]:
@@
The flavor indices $i,j,\ell$ are summed over $N$ values. Imaginary time is $\tau$, we set $\hbar=1$, $s$ is the tuning parameter across the transition, $u$ is a scalar self-interaction. The operator ${\mathcal{D}}_{\bm{r}}=\partial_{x}^{2}-\partial_{y}^{2}$ is special to the Ising-nematic case, and will be set to unity for simplicity in our computations as it unimportant apart from ‘cold spots’ on the Fermi surface. In order to obtain a spin-singlet superconductor we have spinful fermions, with $\sigma$ the spin index. When the couplings are random in flavor space, the model is exactly solved at all $T$ in the large-$N$ limit by (6). Our discussion is for the physical case $N=1$, for which the saddle-point equations in (6) are applicable over an intermediate $T$ regime.
$\mathcal{L}_{\rm 2d-YSYK}$ contains the two sources of disorder: One is the potential $v({\bm{r}})$ acting on the fermions. It is random in both flavor and space (in (2) and (3) we omit flavor indices for clarity, see appendix for full forms):
@@
(i) There is a significant intermediate energy regime where the bosonic and fermionic eigenmodes are spatially extended. The physics is self-averaging, and numerical results are consistent with the large-$N$ averaged Green’s function methods which yield the exact solution of the zero-dimensional YSYK model. This is the regime we address in the present paper by standard SYK methods. The spatial disorder is needed for the singular single-particle self-energies to feed into singular transport properties [8], which does not happen for translationally invariant models [23, 24, 25, 26, 27, 28, 29, 30].
@@
Our numerical findings for the phase diagram are summarized in Fig. 1a, which is plotted as a function of the renormalized boson mass $M$ used to tune the system to the quantum critical point (QCP), and $T$. The phenomenology is broadly similar to that observed experimentally in strange metals: Above the QCP there is a quantum-critical fan, in which the resistivity has an approximately linear $T$ dependence. At low $T$ the QCP is ultimately masked by a superconducting phase, with the maximal $T_{c}$ occurring at the critical point. The d.c. transport is shown in more detail in Fig. 1b. Evidently, the logarithmic corrections to the resistivity mentioned above are relatively weak for the parameters we have considered. On the disordered side of the transition the resistivity is $T$-linear at elevated $T$ inside the quantum critical fan, before going to zero with with an approximately $T^{2}$ power-law below a certain crossover temperature. Over the entire $T$ range shown the resistivity is smaller than the quantum of resistance; the system is not a bad metal.
@@
We now consider the finite-frequency response  [^1] The related problem of optical conductivity at a clean Ising-nematic QCP has been recently analyzed in [57, 58].. The real and imaginary parts of the optical conductivity are presented in Figs. 2a and b. For the detailed structure encoded in $\sigma(\omega)$, we follow Michon et. al. [6] and parametrize $\sigma(\omega)$ via a ‘generalized’ Drude formula:
@@
Here $K$ is the optical weight and is equal to the average electronic kinetic energy (see (LABEL:Kdef) in the supplement [22]), allowing us to determine the frequency-dependent transport scattering rate $1/\tau_{\text{tr}}(\omega)$ and frequency dependent mass-enhancement parameter $m^{*}(\omega)/m$ directly from the data.
@@
The optical scattering rate is shown in Fig. 2c. There is an approximately linear in frequency dependence down to $\omega\sim T$, along with $\omega/T$ scaling (see inset). For frequencies $\omega\lesssim T$, $1/\tau_{\text{tr}}$ tends to a $T$-dependent non-zero value which vanishes in the limit $T\to 0$. At larger $\omega$, the dimensionless ratio $1/[\omega\tau_{\rm tr}(\omega)]$ in Fig. 2c is smaller than observed in [6], but larger values of the ratio appear at larger $g^{\prime}$ (see Fig. LABEL:fig:sig_g_5_insetb [22]).
@@
The frequency-dependent mass enhancement parameter is shown in Fig. 2d. For the chosen parameters, the low-frequency mass enhancement does not exceed roughly 20% down to the lowest temperatures we are able to reliably access in the numerical calculations. The behavior of this “optical” mass enhancement is consistent with the mass enhancement we have inferred from the fermion self-energy [22]. At sufficiently low $T$ $m^{*}(\omega=0)/m$ is expected to diverge logarithmically with $T$ [8]. While we do not observe a pure logarithmic growth of $m^{*}/m$ in our data, the mass enhancement does continuously increase with decreasing $T$ at the QCP, suggesting a (slow) divergence as $T\to 0$.
@@
The modulus and phase of the optical conductivity are shown in Figs. 2e and f: notably, we find the modulus has an apparent sub-linear power law behavior over an intermediate frequency range $|\sigma|\sim 1/\omega^{\nu^{*}}$, where $\nu^{*}\simeq 0.9$ for the chosen set of parameters. Over a similar frequency range the phase of the optical conductivity plateaus at a value $\arg\sigma\simeq\nu^{*}\pi/2$. Such behavior at intermediate frequencies has been observed in infrared conductivity measurements of cuprates [6, 5, 49]. The exponent $\nu^{*}$ is continuously tunable with parameters and, in particular, is a decreasing function of the coupling strength $g^{\prime}$ [22].
@@
In the quantum-critical fan, $\omega/T$ scaling for the optical conductivity is spoiled by logarithmic corrections [8, 6], the most significant such effect arising from the logarithmic divergence of the effective mass with $T$. This may be accounted for by utilizing the generalized Drude formula (4) and considering separately $\omega/T$ scaling of $1/\tau_{\text{tr}}$ and $\Delta m^{*}/m=m^{*}(\omega)/m-m^{*}(0)/m$, where the subtraction removes a contribution expected to violate scaling. With this approach we find reasonable scaling collapse for the optical scattering rate (inset of Fig. 2c). For the parameters presented here, significant logarithmic corrections apparently remain for $\Delta m^{*}/m$. To the extent there is any reasonable scaling collapse, it only holds over a much narrow range of $\omega/T$ (inset of Fig. 2d).
@@
Superconductivity— The superconducting transition temperature $T_{c}$, as shown in Fig. 1a, is numerically identified by the linearized gap equation; see the supplement [22]. For fixed $g^{\prime}$, as we tune away from the QCP by increasing the renormalized boson mass $M$, we find that $T_{c}$ decreases. We compare $T_{c}$ with $A$, the slope of the linear-$T$ resistivity, in the inset of Fig. 1b. The relation is monotonic, but not linear as discussed by Taillefer [19]. We note, however, that in [19] the coefficient $A$ was not extracted in the quantum-critical fan, which is how we have defined it, but rather from fitting to the low-$T$ resistivity in an “extended" critical regime – this regime has been associated with the localization of $\phi$ [9], and is not addressed by our analysis here. If we remain at criticality while changing $g^{\prime}$, then we do find a linear relationship between $A$ and $T_{c}$, as shown in Fig. 6a.
@@
In Fig. 3 we show the evolution of the electronic density of states at the QCP as the system goes through the superconducting transition, demonstrating an apparently conventional gap opening for $T<T_{c}$ despite the fact that, in the absence of superconductivity, the fermions behave as a marginal Fermi liquid with $-\mathrm{Im}\Sigma\sim|\omega|$ at low $T$, as seen in the inset of Fig. 3.
@@
We have also computed the superfluid stiffness $\rho_{s}$ using (8). In Fig. 4, we show the temperature dependence of $\rho_{s}$ for $g^{\prime}=5t^{3/2}$ as the system is tuned away from the QCP by varying the boson mass $M$. The stiffness reaches a finite constant value, $\rho_{s0}$, for $T\rightarrow 0$. As the system is tuned away from the QCP, $T_{c}$ decreases from its maximal value while $\rho_{s0}$ increases, as seen in Fig. 4. This result at strong coupling aligns with the superfluid density in bulk samples of the overdoped cuprate superconductors [50, 20]. When $T_{c}$ is largest at the QCP, the opening of a $\psi$ gap at $T=0$ weakens the pairing interaction mediated by $\phi$ (because the $\phi$ self energy $\Pi$ is self-consistently determined by the $\psi$ polarization, see Fig. 5), leading to a smaller superfluid stiffness at $T=0$ [17, 18]. On the other hand, in the presence of strong potential scattering $v$ this effect is weaker, and a decreasing $T_{c}$ is eventually accompanied by a decreasing $\rho_{s0}$, as shown in Fig. LABEL:fig:stiffness_s [22]. The low $T$ boson localization [9] could have a significant influence on the spatial inhomogeneity of the superfluid density, and this remains to be studied.
@@
Discussion—The spatially inhomogeneous fermion-boson coupling $g^{\prime}({\bm{r}})$ in the 2d-YSYK model in (1) is an alternative to ‘random mass’ $\delta s({\bm{r}})$ disorder in Hertz-Millis theory. The striking similarities between the results presented here and a variety of properties measured in strange metal superconductors indicate such disorder plays a significant role in the phenomenology of these systems, making it important to determine the microscopic origin of disorder configurations that strongly affect the local position of the QCP. The enhancement of weak disorder by strong correlations [51], and the observation of spatial inhomogeneities in correlation-induced ordering [52], are steps in this direction.
@@
Appendix: The 2d-YSYK model and its saddle-point equations – We work with the following lattice action in imaginary time:
$S$ $=\int{\rm d}\tau{\rm d}^{2}x\sum_{i=1}^{N}\sum_{\sigma=\pm 1}\psi^{\dagger}_{i\sigma}(\tau,x)\left[\partial_{\tau}+\varepsilon_{k}\right]\psi_{i\sigma}(\tau,x)+\frac{1}{2}\int{\rm d}\tau{\rm d}^{2}x\sum_{i=1}^{N}\phi_{i}(\tau,x)\left[-\partial_{\tau}^{2}+\omega_{q}^{2}\right]\phi_{i}(\tau,x)$
$+\int{\rm d}^{2}\tau{\rm d}x\sum_{i,j}^{N}\sum_{\sigma=\pm 1}\frac{v_{ij}(x)}{\sqrt{N}}\psi_{i\sigma}^{\dagger}(\tau,x)\psi_{j\sigma}(\tau,x)+\int{\rm d}\tau{\rm d}^{2}x\sum_{i,j,\ell=1}^{N}\sum_{\sigma=\pm 1}\frac{g_{ij\ell}^{{}^{\prime}}(x)}{N}\psi^{\dagger}_{i\sigma}(\tau,x)\psi_{j\sigma}(\tau,x)\phi_{\ell}(\tau,x)\,,$ (1)
where $i,j,\ell$ are flavor indices and $\sigma$ is the spin index. We use the lattice dispersions
$\epsilon_{k}$ $=-2t(\cos k_{x}+\cos k_{y})-\mu\,,$ (2a)
$\omega^{2}_{q}$ $=s+2J(2-\cos q_{x}-\cos q_{y})\,.$ (2b)
Here $t$ is the fermion hopping, $\mu$ is the chemical potential, $s$ is the (squared) bare boson mass and the stiffness $J$ determines the boson dispersion. The Yukawa couplings $g_{ijl}^{{}^{\prime}}$ are complex-valued random variables that obey $g_{ijl}^{{}^{\prime}}(x)=g_{1,ij\ell}(x)+ig_{2,ij\ell}(x)=g_{ji\ell}^{{}^{\prime}}(x)$. The real part $g_{1,ij\ell}(x)$ and the imaginary part $g_{2,ij\ell}(x)$ have zero mean and the variances are [53]
$\overline{g_{1,ij\ell}(x)g_{1,i^{\prime}j^{\prime}\ell^{\prime}}(x^{\prime})}$ $=$ $\left(1-\frac{\alpha}{2}\right)g^{\prime 2}\delta_{\ell,\ell^{\prime}}\left(\delta_{ii^{\prime}}\delta_{jj^{\prime}}+\delta_{ij^{\prime}}\delta_{ji^{\prime}}\right)\delta(x-x^{\prime}),$
$\overline{g_{2,ij\ell}(x)g_{2,i^{\prime}j^{\prime}\ell^{\prime}}(x^{\prime})}$ $=$ $\frac{\alpha}{2}g^{\prime 2}\delta_{\ell,\ell^{\prime}}\left(\delta_{ii^{\prime}}\delta_{jj^{\prime}}-\delta_{ij^{\prime}}\delta_{ji^{\prime}}\right)\delta(x-x^{\prime}),$
$\overline{g_{1,ij\ell}(x^{\prime})g_{2,i^{\prime}j^{\prime},\ell^{\prime}}(x^{\prime})}$ $=$ $0$ (3)
In the $\alpha=1$ limit, (3) reduces to $\overline{g_{ij\ell}^{{}^{\prime}}(x)g_{i^{\prime}j^{\prime}\ell^{\prime}}^{{}^{\prime}}(x^{\prime})^{*}}=g^{\prime 2}\delta(x-x^{\prime})\delta_{ii^{\prime}}\delta_{jj^{\prime}}\delta_{kk^{\prime}}$ and no superconductivity occurs [53]. For $\alpha=0$, however, the coupling constants are all real valued, which preserves time-reversal symmetry for each realization of the random couplings and thus gives rise to superconductivity. However, note that the saddle-point equations in (6) below are independent of $\alpha$ in the normal state. The spatially random potential satisfies
$\overline{v_{ij}(x)}=0\,,\quad\overline{v^{\ast}_{ij}(x)v_{i^{\prime}j^{\prime}}(x^{\prime})}=v^{2}\,\delta(x-x^{\prime})\delta_{ii^{\prime}}\delta_{jj^{\prime}}\,,$ (4)
To tune the system to criticality, we have found it most convenient to first impose a fixed length constraint [33]:
$\sum_{q}\sum_{i=1}^{N}\phi_{iq}(\tau)\phi_{i,-q}(\tau)=N/\gamma\,.$ (5)
This equation implicitly determines $s$, using $\gamma$ as tuning parameter to access the QCP. In the main text we present the phase diagram in terms of the $T=0$ value of the renormalized boson mass $M^{2}=s-\Pi(\omega=0,T=0)$.
After a disorder average, and in the limit of a large number of flavors where the saddle-point approximation is exact, we obtain the following SYK-type equations for the electron Green’s function $\hat{G}$ (a matrix in Nambu space) and the boson Green’s function $D$:
$\hat{G}(i\omega)$ $=$ $\int\frac{{\rm d}^{2}k}{(2\pi)^{2}}\left(i\omega\sigma_{0}+\mu\sigma_{3}-\varepsilon_{k}\sigma_{3}-\hat{\Sigma}(i\omega)\right)^{-1},$ (6a)
$D(i\nu)$ $=$ $\int\frac{{\rm d}^{2}q}{(2\pi)^{2}}\frac{1}{\nu^{2}+\omega_{q}^{2}-\Pi(i\nu)},$ (6b)
$\Sigma(\tau)$ $=$ $g^{\prime 2}G(\tau)D(\tau)+v^{2}G(\tau),$ (6c)
$\Phi(\tau)$ $=$ $-(1-\alpha)(g^{\prime 2}F(\tau)D(\tau)+v^{2}F(\tau)),$ (6d)
$\Pi(\tau)$ $=$ $-2g^{\prime 2}\left(G(\tau)G(-\tau)-(1-\alpha)F(\tau)F^{\dagger}(-\tau)\right),$ (6e)
$\frac{1}{\gamma}$ $=$ $T\sum_{\nu}\int\frac{{\rm d}^{2}q}{(2\pi)^{2}}\frac{1}{\nu^{2}+\omega_{q}^{2}-\Pi(i\nu)}\,.$ (6f)
Here the $\sigma_{i}$ are Pauli matrices in Nambu space, and we use the following parametrization for $\hat{G}$ and $\hat{\Sigma}$
$\hat{G}(i\omega)$ $=$ $G(i\omega)\sigma_{0}+F(i\omega)\sigma_{1},$ (7a)
$\hat{\Sigma}(i\omega)$ $=$ $\Sigma(i\omega)\sigma_{0}+\Phi(i\omega)\sigma_{1}\,.$ (7b)
These equations are indicated schematically in Figure 5 for $v=0$, including the diagrammatic representation of the current-current correlation function used to determine the conductivity $\sigma$. Different from the translationally invariant model, the self-energies are momentum independent as a result of the extra $\delta$-function in (3) and (4).
Appendix: Correlation between $T_{c}$ and the slope of linear in $T$ resistivity – We have also studied the correlation between $A$ – the slope of the $T$ linear resistivity – and $T_{c}$ when the system is tuned precisely to the QCP at $\gamma_{c}$ ($M=0$). In this case $T_{c}$ is tuned by varying the interaction strength $g^{\prime}$. We find an essentially linear relationship, as may be seen in Fig. 6(a). The corresponding tuning parameter $\gamma_{c}$ as a function of $g^{\prime}$ is shown in Fig. 6(b).
Figure 6: (a) Relationship between slope $A$ of the linear in $T$ resistivity and transition temperature $T_{c}$. Here $T_{c}$ is varied by varying $g^{\prime}$ while keeping the system at the QCP. (b) The (inverse) value of the tuning parameter $\gamma$ at the critical point as a function of $g^{\prime}$.
Appendix: Superfluid stiffness – The superfluid stiffness may be obtained from the expectation value of the electronic kinetic energy $K$ and the Matsubara current-current correlation function $\Lambda(i\omega_{n})$ (see (LABEL:eq:SrhoK)) as [56]
$\frac{\rho_{s}}{\pi e^{2}}=\langle-K\rangle/2-\Lambda(i\omega_{n}=0).$ (8)
This result may be expressed purely in terms of the anomalous Green’s function:
$\frac{\rho_{s}}{\pi e^{2}}=2T\sum_{n}\int{\rm d}\epsilon\rho_{\text{tr}}(\epsilon)F^{\dagger}(\epsilon,i\omega_{n})F(\epsilon,i\omega_{n}).$ (9)
The detailed derivation of this result may be found in the supplement.
See pages 1 of supp5 See pages 2 of supp5 See pages 3 of supp5 See pages 4 of supp5 See pages 5 of supp5 See pages 6 of supp5 See pages 7 of supp5 See pages 8 of supp5 See pages 9 of supp5 See pages 10 of supp5 See pages 11 of supp5 See pages 12 of supp5 See pages 13 of supp5 See pages 14 of supp5 See pages 15 of supp5